OpenStudy (anonymous):

What am I doing wrong with this problem? f(x)=3sqrt(5-x)+3 y=3sqrt(5-x)+3 3sqrt(5-x)=y-3 (3sqrt(5-x)^3=(y-3)^3 5-x=(y-3)^3 x=-(y-3)^3-5

OpenStudy (amistre64):

y = 3rt(5-x) + 3 ; i assume you wanna invert it x = (5-y)^(1/3) + 3 x-3 = (5-y)^(1/3) (x-3)^3 = 5-y (x-3)^3 -5 = -y -(x-3)^3 +5 = y

OpenStudy (amistre64):

that last part there; you subtracted the 5; then for got to negate it with the /-1

OpenStudy (amistre64):

[(x-3)^3 -5 = -y] * -1 -((x-3)^3 -5) = y

OpenStudy (anonymous):

I see, ok. thank you.

OpenStudy (amistre64):

youre welcome :) the other day I did 3*3 = 6 twice in a row and couldnt see for the life of me what i did wrong lol

OpenStudy (anonymous):

haha, that kinda week.... Thansk again

OpenStudy (anonymous):

Actually, I just checked this answer and wamap isn't happy with this one either.